# SupplementaryMethods Ultra-high dose rate 6 MeV electron irradiation generates stable [1-13C]alanine radicals suitable for medical imaging with dissolution Dynamic Nuclear Polarisation

### v1

@@ 4.1 Sample Preparation
The samples used in this study can be categorised into three main groups: alanine containing endogenous irradiation-generated radicals, alanine mixed with the exogenous trityl radical OX063 at a concentration of , and pyruvate with AH1111501. These exogeneous radicals differ only by an exterior methyl ‘R’-group in their chemical structure and have almost identical electronic properties; these concentrations of both trityl radicals were chosen to reflect reported optimal conditions for DNP for both molecules.
To prepare the former, dry [1-13C]L-alanine powder (Isotec and Cambridge Isotope Laboratories, Inc) was irradiated using an ultra-high dose rate 6 MeV electron linear accelerator operating in pulsed mode, delivering of dose to the sample in each of the pulses, with a pulse repetition rate of and an approximate beam current of . The dose of irradiation was varied between 10 – depending on the desired radical concentration and was linear with the total number of pulses delivered. The dosimetry and beam delivery was controlled using a beam monitor, which has been described in detail elsewhere,[16, 17] and calibrated against radiochromic film (EBT-XD, Ashland Inc., Covington, KY, USA). The total time for irradiation was deliberately lengthened to avoid linac heating; approximately was irradiated in . The irradiated powder was then used as the main component for two different sample types: either an irradiated dry powder or an irradiated powder/anhydrous glycerol mix with a 1:1 w/w ratio.
The alanine preparation containing OX063 trityl radical (molecular weight = ; GE Healthcare) comprised of [1-13C]L-alanine mixed in 8 ml of water and of HCl. This mixture was freeze-dried, and approximately of alanine hydrochloride yielded. The alanine hydrochloride was then dissolved in DMSO with the assistance of heating. OX063 trityl was added to achieve approximately radical concentration. of this mixture was pipetted into a sample cup.
For the pyruvate sample, [1-13C]pyruvic acid was mixed with the appropriate mass of AH111501 trityl (GE Healthcare; CID 11607875) to achieve a radical concentration, of which was pipetted into a sample cup.
@@ 4.2 Dynamic Nuclear Polarisation
Samples were polarised at 3.35 T or 6.7 T using a HyperSense (Oxford Instruments) or SpinAligner (Polarize), respectively. Further details are provided in the SI §LABEL:si:DNP.
@@ 4.4 In Vivo Experiments
Two healthy male Sprague-Dawley rats were injected with one shot of each type of hyperpolarised alanine sample via a tail vein catheter. The time between each injection was approximately two hours. The respiratory rate and body temperature of each rat were monitored using an MRI compatible small-animal monitoring system (Small Animal Instruments Inc, USA) whilst the rats were under anaesthesia with 2.5-3% sevoflurane in medical air. Both rats were scanned using a dual-tuned 13C/1H volume rat coil at (GE MR750, GE Healthcare). One rat was scanned using a slice-selective spectroscopic acquisition, the other imaged using a 2D MR Spectroscopic Imaging (MRSI) sequence, EPSI, using fidall.
Prior to these scans, a urea phantom was used to calibrate the applied flip angles, and proton images were acquired for anatomical reference. For the slice-selective spectroscopic acquisition, slice thickness was used to ensure the whole abdomen of the rat was imaged. of hyperpolarised sample was injected over approximately 5 s, and the scan started upon the start of the injection. A flip angle was applied with a repetition time (TR) of for a total of 5 minutes (100 time points). For the MRSI, of hyperpolarised substrate was injected over approximately , and the scan started upon the start of the injection. A matrix was chosen, and a FA was applied with a TR of for a total of 2 minutes (8 time points and approximately per time point). The pH of the injected solutions was measured to ensure physiological compatibility, with values ranging from 6.7 to 7.4, well within the acceptable range for intravenous administration. All animal experiments were conducted following appropriate independent ethical review and under appropriate licensing regimes in Denmark.
@@ 4.5 Electron Paramagnetic Resonance
An X-band continuous wave (CW) spectrometer (EMXmicro, Bruker BioSpin) was used to record EPR spectra at room temperature. For recording spectra at and an estimate of , an X-band CW spectrometer (E680, Bruker BioSpin) was used.
Magnesium Oxide powder, a common standard reference in EPR,[72] was used to account for changes in the magnetic field between experiments due to differences in sample positioning.
Samples that comprised dry powder were directly loaded into thin wall quartz EPR tubes, and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25–. Liquid state samples were first loaded in diameter capillary tubes, which in turn were loaded into the EPR tubes.
@@ 4.6 X-ray Diffraction
XRD was conducted using copper X-rays, of wavelength and , in an Empyrean (Malvern Panalytical) diffractometer. XRD samples comprised dry alanine powder or an alanine powder/glycerol mix (1:1 w/w). The Cu wavelength was removed from the diffraction signal with a Ni filter placed in front of the PIXcel1D detector. A set of programmable divergence slits were used for the incident optic, set to illuminate a fixed surface area of across the angular range of the scan. During the scans, the samples were rotated about the azimuthal axis at a frequency of to fully characterise the crystallite distribution.
5 Conflict of Interest
JYCL, CL, JJM, CR and DT are listed as authors on a filed patent (EP24168322.6) pertaining to the use of ultra-high dose rate electron irradiation for preparing hyperpolarised contrast agents.
@@ Acknowledgments
We would like to acknowledge support from the Novo Nordisk Foundation (NNF21OC0068683); the British Heart Foundation (refs. RG/11/9/28921, RE/13/1/30181, FS/19/18/34252, and FS/14/17/30634); the Wellcome Trust (221805/Z/20/Z); the Engineering and Physical Sciences Research Council (EPSRC) and Medical Research Council (MRC) (EP/L016052/1); and the Colleges of St. Hugh’s and Somerville in the University of Oxford. We would furthermore like to acknowledge the University of Oxford British Heart Foundation Centre for Research Excellence (RE/13/1/30181) and the NHS National Institute for Health Research Oxford Biomedical Research Centre programme. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. We also gratefully acknowledge Cambridge Isotope Laboratories, Inc, who provided a gift of [1-13C]alanine to J.J.M. and Z.R. We also thank Christiane Timmel for useful discussions about quantitative EPR spectroscopy. JJM would additionally like to thank Professor Elspeth Garman and Professor Ian Carmichael for both dinner and many useful discussions about radiation chemistry in crystals. The practical assistance of the Department of Physics Mechanical Workshops is also kindly noted.
Author Contributions

### v2

@@ 4.1 Sample Preparation
The samples used in this study can be categorised into three main groups: alanine containing endogenous irradiation-generated radicals, alanine mixed with the exogenous trityl radical OX063 at a concentration of 38 mM, and pyruvate with 30 mM AH1111501. These exogeneous radicals differ only by an exterior methyl ‘R’-group in their chemical structure and have almost identical electronic properties; these concentrations of both trityl radicals were chosen to reflect reported optimal conditions for DNP for both molecules.
To prepare the former, dry [1-13C]L-alanine powder (Isotec and Cambridge Isotope Laboratories, Inc) was irradiated using an ultra-high dose rate 6 MeV electron linear accelerator operating in pulsed mode, delivering 24 Gy of dose to the sample in each of the 4 s pulses, with a pulse repetition rate of 25 Hz and an approximate beam current of 100 mA. Each pulse delivered approximately 480 nC of electric charge. The dose of irradiation was varied between 10 – 100 kGy depending on the desired radical concentration and was linear with the total number of pulses delivered. The dosimetry and beam delivery was controlled using a beam monitor, which has been described in detail elsewhere, (?, ?) and calibrated against radiochromic film (EBT-XD, Ashland Inc., Covington, KY, USA). The total time for irradiation was deliberately lengthened to avoid linac heating; approximately 2 g was irradiated in 160 s. The irradiated powder was then used as the main component for two different sample types: either an irradiated dry powder or an irradiated powder/anhydrous glycerol mix with a 1:1 w/w ratio.
The alanine preparation containing OX063 trityl radical (molecular weight = 1427 g/mol; GE Healthcare) comprised 1.276 g of [1-13C]L-alanine mixed in 8 ml of water and 2 ml of 11.65 M HCl. This mixture was freeze-dried, and approximately 1.5 g of alanine hydrochloride yielded. The alanine hydrochloride was then dissolved in 2.25 ml DMSO with the assistance of heating. 180 mg OX063 trityl was added to achieve approximately 38 mM radical concentration. 50 l of this mixture was pipetted into a sample cup.
For the pyruvate sample, 14 M [1-13C]pyruvic acid was mixed with the appropriate mass of AH111501 trityl (GE Healthcare; CID 11607875) to achieve a 30 mM radical concentration, of which 18 l was pipetted into a sample cup. Dissolutions were performed with either D2O (irradiated alanine) or a D2O solution containing an equimolar amount of NaOH to neutralise.
@@ 4.2 Dynamic Nuclear Polarisation
Samples were polarised at 3.35 T or 6.7 T using a HyperSense (Oxford Instruments) or SpinAligner (Polarize), respectively. Further details are provided in the SI §6.2.
@@ 4.4 In Vivo Experiments
Two healthy male Sprague-Dawley rats were injected with one shot of each type of hyperpolarised alanine sample via a tail vein catheter. The time between each injection was approximately two hours. The respiratory rate and body temperature of each rat were monitored using an MRI compatible small-animal monitoring system (Small Animal Instruments Inc, USA) whilst the rats were under anaesthesia with 2.5-3% sevoflurane in 1 l/min medical air. Both rats were scanned using a dual-tuned 13C/1H volume rat coil at 3 T (GE MR750, GE Healthcare). One rat was scanned using a slice-selective spectroscopic acquisition, the other imaged using a 2D MR Spectroscopic Imaging (MRSI) sequence, EPSI, using fidall. (?)
Prior to these scans, a urea phantom was used to calibrate the applied flip angles, and proton images were acquired for anatomical reference. For the slice-selective spectroscopic acquisition, 80 mm slice thickness was used to ensure the whole abdomen of the rat was imaged. 1 ml of hyperpolarised sample was injected over approximately 5 s, and the scan started upon the start of the injection. A 30 ∘ flip angle was applied with a repetition time (TR) of 3 s for a total of 5 minutes (100 time points). For the MRSI, 1 ml of hyperpolarised substrate was injected over approximately 5 s, and the scan started upon the start of the injection. A matrix was chosen, and a 15 ∘ flip angle was applied with a TR of 64 ms for a total of 2 minutes (8 time points and approximately 16 s per time point). The pH of the injected solutions was measured to ensure physiological compatibility, with values ranging from 6.7 to 7.4, well within the acceptable range for intravenous administration. All animal experiments were conducted following appropriate independent ethical review and under appropriate licensing regimes in Denmark.
@@ 4.5 Electron Paramagnetic Resonance
An X-band continuous wave (CW) spectrometer (EMXmicro, Bruker BioSpin; Super High Sensitivity rectangular cavity probe head) was used to record EPR spectra at room temperature. For recording spectra at 5 K and measurement of and , an X-band pulsed spectrometer (E680, Bruker BioSpin) was used. The Chromium(III) impurity in Magnesium Oxide powder, a common standard in EPR, (?, ?) was used as a frequency reference. Further details are provided in SI §6.7.
Samples that comprised dry powder were directly loaded into 4 mm thin wall quartz EPR tubes, and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25–35 ∘. Liquid state samples were first loaded in 2 mm diameter capillary tubes, which in turn were loaded into the EPR tubes.
@@
13C spectra were fitted via a constrained Bayesian optimisation routine (?). The line widths and shapes of spectra were also compared to models built with the ‘pepper’ function in EasySpin, a validated QM spectral simulation library designed for EPR. (?)
@@ 4.6 X-ray Diffraction
XRD was conducted using copper X-rays, of wavelength 1.5405982 Å and 1.544497 Å, in an Empyrean (Malvern Panalytical) diffractometer. XRD samples comprised dry alanine powder or an alanine powder/glycerol mix (1:1 w/w). The Cu wavelength was removed from the diffraction signal with a Ni filter placed in front of the PIXcel1D detector. A set of programmable divergence slits were used for the incident optic, set to illuminate a fixed surface area of 1 cm2 across the angular range of the scan. During the scans, the samples were rotated about the azimuthal axis at a frequency of 0.5 Hz to fully characterise the crystallite distribution.
XRD data were compared to a previously reported crystal structure of alanine, obtained via neutron scattering (?), and a Rietveld refinement was undertaken in Maud using a spherical harmonic texture basis (?). Further details are provided in the SI §6.9. [7890 Words]
5 Figures and Tables
Fig. | Sample/Condition | Dose | Build-up | Fast | Slow | |
(kGy) | Amplitude | Time | Time | Weighting | ||
(a.u.) | (s) | (h) | ||||
2a | 10 kGy (glycerol plug) | 10 | Monoexponential fit poor; s | |||
100 kGy (glycerol plug) | 100 | ( s) | ||||
100 kGy (glycerol mix) | 100 | ( s) | ||||
3b | 30 kGy (glycerol mix) | 30 | ||||
50 kGy (glycerol mix) | 50 | |||||
70 kGy (glycerol mix) | 70 | |||||
4c | 13C-Pyruvate + OX063 | – | ( s) | |||
13C-Alanine + OX063 | – | ( s) | ||||
70 kGy (glycerol mix) | 70 |
Notes: These data are at 3.35 T.
OX063 trityl radical shows purely monoexponential behaviour ( in one fit; identical time constants in the other).
@@ Acknowledgments
We would like to acknowledge support from the Novo Nordisk Foundation (NNF21OC0068683); the British Heart Foundation (refs. RG/11/9/28921, RE/13/1/30181, FS/19/18/34252, and FS/14/17/30634); the Wellcome Trust (221805/Z/20/Z); the Engineering and Physical Sciences Research Council (EPSRC) and Medical Research Council (MRC) (EP/L016052/1); and the Colleges of St. Hugh’s and Somerville in the University of Oxford. We would furthermore like to acknowledge the University of Oxford British Heart Foundation Centre for Research Excellence (RE/13/1/30181) and the NHS National Institute for Health Research Oxford Biomedical Research Centre programme. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. AMB is grateful to The Royal Society and the EPSRC for a Dorothy Hodgkin Fellowship (DH160004 and DHF/R/221018), and the University of Manchester for a Dame Kathleen Ollerenshaw Fellowship. AMB also thanks the Royal Society of Chemistry for a Community for Analytical and Measurement Science fellowship (CAMS Fellowship 2020 ACTF ref 600310/09). AMB is grateful for support from National Research Facility for EPR (EP/W014521/1, NS/A000055/1, EP/V035231/1, and EP/S033181/1). We also gratefully acknowledge Cambridge Isotope Laboratories, Inc, who provided a gift of [1-13C]alanine to JJM and Z.R. We also thank Christiane Timmel for useful discussions about quantitative EPR spectroscopy. JJM would additionally like to thank Professor Elspeth Garman and Professor Ian Carmichael for both dinner and many useful discussions about radiation chemistry in crystals. The practical assistance of the University of Oxford Department of Physics Mechanical Workshops is also kindly noted.
Author contributions: